40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m elog |
Not logged in |
 |
|
Wed Sep 29 17:10:09 2021, Anchal, Summary, CDS, c1teststand problems summary
|
Thu Sep 30 14:09:37 2021, Anchal, Summary, CDS, New way to ssh into c1teststand
|
Thu Mar 3 15:37:40 2022, Anchal, Summary, CDS, c1teststand restructured
|
Mon Oct 4 11:05:44 2021, Anchal, Summary, CDS, c1teststand problems summary
|
Mon Oct 4 18:00:16 2021, Koji, Summary, CDS, c1teststand problems summary
|
Tue Oct 5 17:58:52 2021, Anchal, Summary, CDS, c1teststand problems summary
|
Tue Oct 5 18:00:53 2021, Anchal, Summary, CDS, c1teststand time synchronization working now
|
Mon Oct 11 17:31:25 2021, Anchal, Summary, CDS, Fixed mounting of mx devices in fb. daqd_dc is running now.
|
Mon Oct 11 18:29:35 2021, Anchal, Summary, CDS, Moving forward?
|
Tue Oct 12 17:20:12 2021, Anchal, Summary, CDS, Connected c1sus2 to martian network
|
Tue Oct 12 23:42:56 2021, Koji, Summary, CDS, Connected c1sus2 to martian network
|
Wed Oct 13 11:25:14 2021, Anchal, Summary, CDS, Ran c1sus2 models in martian CDS. All good!
|
Tue Oct 19 18:20:33 2021, Ian MacMillan, Summary, CDS, c1sus2 DAC to ADC test 
|
Tue Oct 19 23:43:09 2021, Koji, Summary, CDS, c1sus2 DAC to ADC test   
|
Wed Oct 20 11:48:27 2021, Anchal, Summary, CDS, Power supple configured correctly.
|
Tue Oct 26 18:24:00 2021, Ian MacMillan, Summary, CDS, c1sus2 DAC to ADC test 
|
Wed Dec 22 17:40:22 2021, Anchal, Summary, CDS, c1su2 model updated with SUS damping blocks for 7 SOSs
|
Wed Dec 29 20:09:40 2021, rana, Summary, CDS, c1su2 model updated with SUS damping blocks for 7 SOSs
|
Fri Mar 4 11:04:34 2022, Anchal, Summary, CDS, c1susaux2 system setup and running
|
Mon Mar 7 19:38:47 2022, Anchal, Summary, CDS, c1susaux2 slow controls issues
|
Mon Mar 14 12:20:05 2022, Anchal, Summary, CDS, c1susaux2 slow controls acromag chassis installed 
|
Thu Mar 17 19:12:44 2022, Anchal, Summary, CDS, c1auxey1 slow controls acromag chassis installed, not powered
|
Fri Mar 18 18:39:13 2022, Yehonathan, Summary, CDS, c1auxey1 slow controls acromag chassis installed, powered
|
Fri Mar 18 19:10:51 2022, Anchal, Summary, CDS, c1auxey1 slow controls issues
|
Mon Mar 21 18:42:06 2022, Anchal, Summary, CDS, c1auxey1 slow controls issues
|
Mon Apr 4 17:03:47 2022, Anchal, Summary, CDS, c1susaux2 slow controls acromag chassis fixed and installed
|
Wed Jul 6 22:40:03 2022, Tega, Summary, CDS, Use osem variance to turn off SUS damping instead of coil outputs
|
Thu Jul 7 21:25:48 2022, Tega, Summary, CDS, Use osem variance to turn off SUS damping instead of coil outputs
|
Tue Mar 15 11:52:34 2022, Anchal, Summary, CDS, c1su2 model updated for sending Run/Acquire Binary Output to Binary Interface card
|
Tue Mar 15 14:10:41 2022, Anchal, Summary, CDS, c1su2 model remade, reinstalled, restarted after the update
|
Tue Oct 12 17:10:56 2021, Anchal, Summary, CDS, Some more information
|
|
Message ID: 16976
Entry time: Wed Jul 6 22:40:03 2022
In reply to: 16756
Reply to this: 16979
|
Author: |
Tega |
Type: |
Summary |
Category: |
CDS |
Subject: |
Use osem variance to turn off SUS damping instead of coil outputs |
|
|
I updated the database files for the 7 BHD optics to separate the OSEM variance trigger and the LATCH_OFF trigger operations so that an OSEM variance value exceeding the max of say 200 cnts turns off the damping loop whereas pressing the LATCH_OFF button cuts power to the coil. I restarted the modbusIOC service on c1susaux2 and checked that the new functionality is behaving as expected. So far so good.
TODO
Figure out the next layer of watchdogging needed for the BHD optics.
Quote: |
[Anchal, JC, Ian, Paco]
We have now fixed all issues with the PD mons of c1susaux2 chassis. The slow channels are now reading same values as the fast channels and there is no arbitrary offset. The binary channels are all working now except for LO2 UL which keeps showing ENABLE OFF. This was an issue earlier on LO1 UR and it magically disappeared and now is on LO2. I think the optical isolators aren't very robust. But anyways, now our watchdog system is fully functional for all BHD suspended optics.
|
|